OpenStudy (anonymous):

solve the given equation log5 (x+2) - log5 11= log5 121 ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use 2nd rule listed there then simplify and cross multiply

OpenStudy (anonymous):

can you show me @jdoe0001

OpenStudy (jdoe0001):

well, let us try using the 2nd rule listed there on the left hand side.... what would that give you?

OpenStudy (jdoe0001):

\(\bf log_5(x+2) - log_5(11)= log_5(121)\implies log_5\left(\cfrac{x+2}{11}\right)= log_5(121) \\ \quad \\ \cfrac{x+2}{11}=121\)

OpenStudy (anonymous):

so is that the answer @jdoe0001

OpenStudy (jdoe0001):

you need to solve for "x"

OpenStudy (anonymous):

how would we do that @jdoe0001

OpenStudy (anonymous):

@jdoe0001 ??

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=n3FbQGGKVfc <---- simplifying linear equations

OpenStudy (anonymous):

@surjithayer ??

OpenStudy (anonymous):

x+2=121*11=? find x

OpenStudy (bmorse):

now multiply: 121*11= x+2 1331= x+2 x= 1331-2 = 1329

OpenStudy (anonymous):

are you sure this is correct @bmorse

OpenStudy (bmorse):

hang on let me check...

OpenStudy (bmorse):

Yes it is!!

OpenStudy (anonymous):

thanks

OpenStudy (bmorse):

you're welcome!